OpenStudy (lilai3):

A standard deck of playing cards contains 52 cards in 4 suits of 13 cards each. 2 suits are read and 2 suits are black. Find each probability. Assume the first card is replaced before the second card is drawn. P(black, queen)

OpenStudy (anonymous):

4. P(black, queen) = 26/52 * 4/52 = 0.0385 5. P(black, diamond) = 26/52 * 13/52 = 0.1250 6. P(jack, queen) = 4/52 * 4/52 = 0.0059 hope this helps

OpenStudy (lilai3):

thank you very much!

OpenStudy (lilai3):

I looked at this and now I just remembered what I thought in my mind when I first came here and you replied all in one time... Here's the question: How did you know I was asking for these answers? Bonus question (I would LOVE you to answer this one: Did and were you sTaLkInG me?